# Complete Synchronization and Pattern Selection through Amplitude Dynamics and Diffusion in Heterogeneous Oscillatory Media

### v1

@@ 1 Introduction
Oscillations are a fundamental dynamical state that determines the functioning of many natural and human-made systems. Macroscopic oscillations are often observed in systems comprising many interacting microscopic oscillatory units and result from synchronization, where the rhythms of the individual oscillating entities are adjusted to a common frequency [1]. This underscores the importance of understanding the emergence of collective behavior in ensembles of coupled oscillators.
@@
Key synchronization features are captured by the paradigmatic Kuramoto model that reduces the dynamics of an oscillator to the dynamics of its phase [2, 3, 4]. In particular, the model captures the transition from an incoherent to a partially synchronized state in an ensemble of globally coupled oscillators with a unimodal frequency distribution as the coupling strength increases. The theoretically predicted second-order-type transition could be validated experimentally with an array of globally coupled electrochemical oscillators [5]. Other diverse experimental realizations range from the synchronization of generic mechanical metronomes [6] to the synchronization of oscillators as complex as Josephson junctions [7].
@@
Because of its simplicity and universality, the Kuramoto model has been extended in various directions. In particular, various coupling topologies, coupling functions, types of frequency distributions, or additional external fields have been employed. Comprehensive reviews summarize these extensions and some of their applications in different disciplines [8, 9].
@@
However, there exist synchronization patterns that exhibit pronounced amplitude variations [10, 11, 12]. This typically happens when the coupling strength is large. Clearly, phase models fail in this coupling regime. Instead, modeling approaches often use networks of Stuart-Landau (SL) oscillators, generic oscillators close to a supercritical Hopf bifurcation [13]. Compared to phase models, the complexity of systems of coupled SL oscillators is significantly greater, and our understanding of their dynamics remains in its infancy. Most investigations consider identical and globally coupled SL oscillators with a focus on identifying qualitatively new types of collective states. Among the described collective dynamics are amplitude cluster solutions, chimera states, and incoherent solutions that adapt amplitudes off the unit circle [14, 15, 16, 17, 18]. Other studies investigate the multiplicity of the solutions which result from the inherent full permutation symmetry of systems of identical oscillators [19, 20, 21, 22, 23]. In contrast, the literature on the impact of heterogeneous frequency distributions in an ensemble of SL oscillators is still sparse and focused on a few parameter regimes [24, 25, 26].
@@
So far, we have considered ensembles of discrete oscillators. In many physical systems, however, the situation is more involved. This applies in particular to globally coupled oscillating continuous media, where, in addition to the global coupling, oscillators interact with each other due to transport processes such as diffusion. Prominent examples are systems that are electrically controlled, such as gas discharge tubes [27] and electrode reactions [28], but also the light-sensitive Belousov-Zhabotinsky reaction [29, 30], or the heterogeneous CO oxidation on Pt single crystal catalysts [31]; observed patterns include cluster formation, standing waves, and stripe pattern. The theoretical description of these systems was achieved with the complex Ginzburg-Landau equation (CGLE) extended by a global coupling term [31, 32].
@@
Here, we introduce a type of amplitude pattern that forms in experiments involving an oscillatory continuous medium with a strong global and a weak diffusional coupling. The pattern is captured by an extended CGLE with small heterogeneities. However, since it is dominated by global coupling, the mechanism by which it forms can be elucidated using a heterogeneous system of globally coupled SL oscillators. Furthermore, we demonstrate that the solution inherits its ordering principle in parameter space from the cluster singularity [20] found in globally coupled identical SL oscillators. In addition, treating diffusional coupling as a small perturbation, we can derive a Lyapunov functional and demonstrate that in the presence of a small diffusion, the system selects a particular state among the large amount of coexisting stable patterns in discrete globally coupled oscillators. We thus combine the various approaches mentioned above to understand collective oscillations, enabling us to demonstrate how disturbances in real systems influence their dynamics.
@@
Experimentally, we study the anodic electrochemical dissolution of silicon, a wet-chemical preparation method for obtaining smooth Si surfaces [33]. When silicon is immersed in fluoride-containing electrolyte under anodic bias, the dissolution process does not proceed steadily, but instead generates self-sustaining oscillations in oxide thickness and current [34, 35]. These oscillations indicate nonlinear feedback loops between interfacial reactions, charge transport, and oxide growth. When using n-doped Si wafers, patterns in the oxide layer thickness are observed where typically the amplitudes of different oscillating regions vary significantly  [36, 11]. In this article, we introduce a particular experimentally obtained collective state: a bimodal amplitude solution, i.e., a pattern characterized by two regions, in which the amplitudes of the oscillators are centered around one of the two modes within a region. At the same time, the entire electrode oscillates at a uniform frequency, i.e., it exhibits complete frequency synchronization. Interestingly, the ratio between the two amplitude areas could be continuously tuned via a global coupling parameter.
@@ 2 Experiment with Electrochemical Silicon Dissolution System
The impact of global coupling on the formation of amplitude patterns was investigated experimentally using an electrochemical oscillator, silicon electrooxidation in a fluoride-containing electrolyte. For an overview of the anodic behavior of Si electrodes, see chapter 5 in [33], details of the oscillation mechanism can be found in [37]. The Si oscillator is known to exhibit amplitude clusters at certain parameter values [12]. In this study, we control the degree of global coupling and demonstrate how it influences the characteristics of the amplitude patterns.
@@ 2.1 Setup
Illumination is performed with a laser (He–Ne, $632\,\rm nm$) to generate valence band holes necessary for the electrochemical reaction. The illumination intensity is adjusted to $0.88\,\rm mW/cm^{2}$ so amplitude clusters can be observed [12] using an electrically controlled gray filter.
During the electrochemical measurements, an ellipsometric imaging device monitors the spatiotemporal evolution of the silicon oxide layer formed at the interface between the working electrode and the electrolyte. The recorded signal is the light intensity reflected at the interface. The signal varies with the optical path length, including changes in the oxide’s thickness and the dielectric constant at the interface. The details of the ellipsometric setup are described in Ref. [38].
@@
where the integration runs over the electrode area, $A$. The electrode area was $A=14.52\,\rm mm^{2}$ in all measurements. As the local current is a function of $U_{\rm eff}$, the external resistor induces a global coupling. The higher the resistance, the stronger the global coupling. We use $R=R_{\rm ext}\cdot A\,\rm[k\Omega cm^{2}]$ to quantify the global coupling intensity hereafter. To keep the local reaction rate practically independent of the strength of the global coupling, we adjusted $U_{\rm app}$ such that $U_{\rm eff}$ was always in the same range. More precisely, $R_{\rm ext}$ was changed in $3\,\rm k\Omega$ increments and at each step $U_{\rm app}$ was increased by $0.13\,\rm V$. The following table shows the pairs of investigated parameters $R$ and $U_{\rm app}$.
@@ 2.2 Data analysis
where the kernel $K$ is a Gaussian. The bandwidth $h$ controls the smoothing; its default value was chosen according to Silverman’s rule of thumb [39].
Figure 1: Experimentally measured distribution of time-averaged amplitude (first row), snapshots of phase distribution (second row), and time-averaged frequency distribution (third row) on an electrode surface during anodic Si electrodissolution. From (a) to (d) the external resistance was increased: (a) $R=0.58\rm\,k\Omega cm^{2}$, (b) $R=1.46\rm\,k\Omega cm^{2}$, (c) $R=2.33\rm\,k\Omega cm^{2}$, and (d) $R=3.65\rm\,k\Omega cm^{2}$.
@@ 2.3 Results
Figure 3: Relative size $\rho$ of the experimentally obtained high-amplitude region as a function of the external resistance $R$.
@@ 3.1 Modified Complex Ginzburg-Landau Equation
To model the experiments, we turn to a modified complex Ginzburg–Landau equation (CGLE), which is commonly used in the study of reaction-diffusion systems [3, 40]. The CGLE naturally captures the local oscillatory dynamics via Stuart–Landau terms, which describe the silicon system that undergoes a Hopf bifurcation, as described in Ref. [37]. We add a mean-field coupling term to incorporate the experimentally observed control of the ratio of high- and low-amplitude regions through the external resistance. This accounts for the global electrical feedback (see Section 2) acting equally on all oscillators. Since perfectly homogeneous conditions cannot be achieved in any electrochemical experiments, we consider spatial heterogeneity by assuming that the intrinsic frequencies $\omega(\vec{x})$ vary slightly across the electrode surface. Finally, spatial coupling through ionic and electronic transport must be considered because the silicon electrode is an extended medium; this effect is modeled by a diffusion term. Taken together, we end up with the heterogeneous, globally coupled CGLE:
@@
Here, $\vec{x}=(x_{1},x_{2})$ denotes the spatial coordinates on the rectangular domain $0,L_{1}]\times[0,L_{2}]$. The global coupling parameters are $K$ and $C_{1}$, where $K$ sets the coupling strength and $C_{1}$ modulates the phase shift introduced by the coupling. The shear parameter of the individual oscillators, $C_{2}$, quantifies the amplitude–frequency coupling. The diffusion strength is given by $\epsilon$. The intrinsic frequencies $\omega_{k}$ are drawn from a Gaussian distribution with mean $\mu_{0}$ and standard deviation $\sigma_{0}$; moving to a co-rotating frame with $\mu_{0}$ as its center shifts the distribution to zero.
Figure 4: Numerically obtained snapshots of the amplitude (first row) and phase (second row) distributions, and time-averaged frequency distribution (third row) for $K=1.15,C_{2}=2,\epsilon=0.02,\sigma_{0}=0.3$. First column $C_{1}=-2.5$, second column $C_{1}=-2.1$, third column $C_{1}=-1.6$, and last column $C_{1}=-1.4$.
We numerically solve Eqs. (5) on a square domain ($L_{1}=L_{2}=10$) with Neumann boundary conditions, discretized on a $100\times 100$ uniform grid. To ensure reproducible frequency realizations, we generate $\omega(\vec{x})$ with Julia’s Mersenne Twister PRNG (seed 2024). We use adiabatic continuation and integrate system (5) for 1000 time steps. The numerical integrations are performed using Julia’s DifferentialEquations.jl ecosystem [41]. We use Tsit5 [42], a Runge-Kutta-based solver with adaptive timesteps, and save data points at $\Delta t=0.1$.
Starting from $C_{1}=-1.75$, we vary the coupling parameter in increments of $\pm 0.05$ until reaching $C_{1}=-1$ and $C_{1}=-3$, respectively. Figure 4 shows the numerical results of the spatial distribution of the amplitude (first row), phase (second row), and frequency (third row) for four different $C_{1}$-values. The amplitude and phase distributions are shown as snapshots, while the frequency is averaged over the last 50 integration steps. All results are recorded after a sufficient evolution time to ensure that transients have decayed.
@@
In Figure 5, we plotted the distribution of the amplitudes for the same $C_{1}$-values used in Figure 4. From the data, it is clear that there are two peaks for all four values of $C_{1}$. Using the Kernel density estimate defined in equation (4), we can locate the mean radius of each mode. The maxima of the distributions, shown as blue dots, are used to quantify the mode states via the ratio $\rho$ defined in Section 2. Since the experimental peaks are roughly twice as broad as the simulated ones, we restrict the counting to oscillators within $\pm 0.1$ of each peak maximum. From the four different histograms, one can clearly observe how oscillators transition from the low-amplitude mode to the high-amplitude mode as $C_{1}$ is increased.
@@
In Figure 6, we plotted the fraction of the high-amplitude mode $\rho$ as a function of $C_{1}$ for three different simulation schemes. A unimodal distribution is assigned to $\rho=0$. The blue open circles represent simulation results, where, for $C_{1}=-1.75$, a random initial condition was chosen. For the other values of $C_{1}$, the simulations were continued adiabatically. For the simulations shown in orange symbols, the same initial condition is chosen for each value of $C_{1}$, and the simulation shown in green is identical to orange, but without diffusion. First, consider the adiabatic continuation. Starting at $C_{1}=-3$, the system is in a synchronous solution. At $C_{1}\approx-2.8$, it jumps to a pattern of $\rho\approx 0.06$. In this configuration, many oscillators are in the low-amplitude region, and few are in the high-amplitude region. Increasing $C_{1}$ leads to more oscillators transitioning to the high amplitude region until a critical ratio is reached at $C_{1}\approx-1.1$, at which the system transitions back to a unimodal distribution. For the continuation with a random initial condition, the $C_{1}$-interval where the bimodal pattern appears is narrower. For the no-diffusion series (green), the amplitude distribution is often not clearly bimodal; in those cases, the KDE peak–picking ansatz becomes ill-posed and the inferred $\rho$ fluctuates strongly. Therefore, these apparent “jumps” should be interpreted cautiously, as they largely reflect estimator instability rather than genuine rapid changes of the domain fraction.

### v2

# Synchronized Bimodal Amplitude Patterns in Heterogeneous Oscillatory Media — Experiment and Theory

@@ 1 Introduction
Oscillations are a fundamental dynamical state that determines the functioning of many natural and human-made systems. Macroscopic oscillations are often observed in systems comprising many interacting microscopic oscillatory units and result from synchronization, where the rhythms of the individual oscillating entities are adjusted to a common frequency [Pikovsky.2010]. This underscores the importance of understanding the emergence of collective behavior in ensembles of coupled oscillators.
@@
Key synchronization features are captured by the paradigmatic Kuramoto model that reduces the dynamics of an oscillator to the dynamics of its phase [Kuramoto.1975, Haken.1984, Strogatz.2000]. In particular, the model captures the transition from an incoherent to a partially synchronized state in an ensemble of globally coupled oscillators with a unimodal frequency distribution as the coupling strength increases. The theoretically predicted second-order-type transition could be validated experimentally with an array of globally coupled electrochemical oscillators [Kiss.2002]. Other diverse experimental realizations range from the synchronization of generic mechanical metronomes [Pantaleone.2002] to the synchronization of oscillators as complex as Josephson junctions [Wiesenfeld.1998].
@@
Because of its simplicity and universality, the Kuramoto model has been extended in various directions. In particular, various coupling topologies, coupling functions, types of frequency distributions, or additional external fields have been employed. Comprehensive reviews summarize these extensions and some of their applications in different disciplines [Acebron.2005, Gupta.2014].
@@
However, there exist synchronization patterns that exhibit pronounced amplitude variations [Wang.2000, Schmidt.2014, Patzauer.2021]. This typically happens when the coupling strength is large. Clearly, phase models fail in this coupling regime. Instead, modeling approaches often use networks of Stuart-Landau (SL) oscillators, generic oscillators close to a supercritical Hopf bifurcation [Guckenheimer.1983]. Compared to phase models, the complexity of systems of coupled SL oscillators is significantly greater, and our understanding of their dynamics remains in its infancy. Most investigations consider identical and globally coupled SL oscillators with a focus on identifying qualitatively new types of collective states. Among the described collective dynamics are amplitude cluster solutions, chimera states, and incoherent solutions that adapt amplitudes off the unit circle [Nakagawa.1993, Nakagawa.1994, Chabanol.1997, Sethia.2014, Clusella.2019]. Other studies investigate the multiplicity of the solutions which result from the inherent full permutation symmetry of systems of identical oscillators [Golubitsky.2002, Kemeth.2019, Kemeth.2020, Ku.2015, Thome.2025]. In contrast, the literature on the impact of heterogeneous frequency distributions in an ensemble of SL oscillators is still sparse and focused on a few parameter regimes [Aronson.1990, Matthews.1991, Monte.2002, Millan.2025].
@@
So far, we have considered ensembles of discrete oscillators. In many physical systems, however, the situation is more involved. This applies in particular to globally coupled oscillating continuous media, where, in addition to the global coupling, oscillators interact with each other due to transport processes such as diffusion. Prominent examples are systems that are electrically controlled, such as gas discharge tubes [Ammelt.1997] and electrode reactions [Krischer.2002], but also the light-sensitive Belousov-Zhabotinsky reaction [Vanag.2000, Nicola.2006], or the heterogeneous CO oxidation on Pt single crystal catalysts [Falcke.1995]; observed patterns include cluster formation, standing waves, and stripe pattern. The theoretical description of these systems was achieved with the complex Ginzburg-Landau equation (CGLE) extended by a global coupling term [Falcke.1995, Yochelis.2004].
@@
Here, we introduce a type of amplitude pattern that forms in experiments involving an oscillatory continuous medium with a strong global and a weak diffusional coupling. The pattern is captured by an extended CGLE with small heterogeneities. However, since it is dominated by global coupling, the mechanism by which it forms can be elucidated using a heterogeneous system of globally coupled SL oscillators. Furthermore, we demonstrate that the solution inherits its ordering principle in parameter space from the cluster singularity [Kemeth.2019] found in globally coupled identical SL oscillators. In addition, treating diffusional coupling as a small perturbation, we can derive a Lyapunov functional and demonstrate that in the presence of a small diffusion, the system selects a particular state among the large amount of coexisting stable patterns in discrete globally coupled oscillators. We thus combine the various approaches mentioned above to understand collective oscillations, enabling us to demonstrate how disturbances in real systems influence their dynamics.
@@
Experimentally, we study the anodic electrochemical dissolution of silicon, a wet-chemical preparation method for obtaining smooth Si surfaces [XiaogeGregoryZhang.2004]. When silicon is immersed in fluoride-containing electrolyte under anodic bias, the dissolution process does not proceed steadily, but instead generates self-sustaining oscillations in oxide thickness and current [Turner.1958, Blackwood.1992]. These oscillations indicate nonlinear feedback loops between interfacial reactions, charge transport, and oxide growth. When using n-doped Si wafers, patterns in the oxide layer thickness are observed where typically the amplitudes of different oscillating regions vary significantly  [Schonleber.2014, Schmidt.2014]. In this article, we introduce a particular experimentally obtained collective state: a bimodal amplitude solution, i.e., a pattern characterized by two regions, in which the amplitudes of the oscillators are centered around one of the two modes within a region. At the same time, the entire electrode oscillates at a uniform frequency, i.e., it exhibits complete frequency synchronization. Interestingly, the ratio between the two amplitude areas could be continuously tuned via a global coupling parameter.
@@ 2 Experiment with Electrochemical Silicon Dissolution System
The impact of global coupling on the formation of amplitude patterns was investigated experimentally using an electrochemical oscillator, silicon electrooxidation in a fluoride-containing electrolyte. For an overview of the anodic behavior of Si electrodes, see chapter 5 in [XiaogeGregoryZhang.2004], details of the oscillation mechanism can be found in [Murakami.]. The Si oscillator is known to exhibit amplitude clusters at certain parameter values [Patzauer.2021]. In this study, we control the degree of global coupling and demonstrate how it influences the characteristics of the amplitude patterns.
@@ 2.1 Setup
Illumination is performed with a laser (He–Ne, $632\,\rm nm$) to generate valence band holes necessary for the electrochemical reaction. The illumination intensity is adjusted to $0.88\,\rm mW/cm^{2}$ so amplitude clusters can be observed [Patzauer.2021] using an electrically controlled gray filter.
During the electrochemical measurements, an ellipsometric imaging device monitors the spatiotemporal evolution of the silicon oxide layer formed at the interface between the working electrode and the electrolyte. The recorded signal is the light intensity reflected at the interface. The signal varies with the optical path length, including changes in the oxide’s thickness and the dielectric constant at the interface. The details of the ellipsometric setup are described in Ref. [Miethe.2009].
@@
where the integration runs over the electrode area, $A$. The electrode area was $A=14.52\,\rm mm^{2}$ in all measurements. In this system, the local electrochemical reaction is mediated by holes in the valence band, and the formation of patterns is known to be governed by the horizontal motion of the holes on the surface of the electrode, resulting from a reaction–diffusion–migration equation with an additional nonlinear, nonlocal term (see Eq. (5) in the Supplementary Material of Ref. [Patzauer.2021]). Therefore, the local density of the holes, controlled by $U_{\rm eff}$, is globally coupled through the external resistor as indicated in equation (1). The higher the resistance, the stronger the global coupling. We use $R=R_{\rm ext}\cdot A\,\rm[k\Omega cm^{2}]$ to quantify the global coupling intensity hereafter. To keep the local reaction rate practically independent of the strength of the global coupling, we adjusted $U_{\rm app}$ such that $U_{\rm eff}$ was always in the same range. More precisely, $R_{\rm ext}$ was changed in $3\,\rm k\Omega$ increments and at each step $U_{\rm app}$ was increased by $0.13\,\rm V$. In each step, the parameter was kept constant for 400 s after the initial transient period. The following table shows the pairs of investigated parameters $R$ and $U_{\rm app}$.
@@ 2.2 Data analysis
where the kernel $K$ is a Gaussian. The bandwidth $h$ controls the smoothing; its default value was chosen according to Silverman’s rule of thumb [Sheather.2004].
Figure 1: Experimentally measured distribution of time-averaged amplitude (first row), snapshots of phase distribution (second row), and time-averaged frequency distribution (third row) on an electrode surface during anodic Si electrodissolution. From (a) to (d) the external resistance was increased: (a) $R=0.58\rm\,k\Omega cm^{2}$, (b) $R=1.46\rm\,k\Omega cm^{2}$, (c) $R=2.33\rm\,k\Omega cm^{2}$, and (d) $R=3.65\rm\,k\Omega cm^{2}$. The resistance was increased in steps, as shown in Table 1. At each step, we waited 400 s after the initial transient state.
@@ 2.3 Results
Figure 3: Relative size of the experimentally obtained high-amplitude region as a function of the resistance.
@@ 3.1 Modified Complex Ginzburg-Landau Equation
To model the experiments, we employ a modified complex Ginzburg–Landau equation (CGLE), a standard amplitude equation for oscillatory reaction–diffusion systems [Haken.1984, GarciaMorales.2012]. The CGLE naturally captures the local oscillatory dynamics through Stuart–Landau terms, which describe the oscillations in the silicon system that arise through a Hopf bifurcation, as explained in Ref. [Schonleber.2014] and more recently justified mathematically in Ref. [Murakami.]. We introduce a general mean-field coupling term to account for the experimentally observed control of the ratio of high- and low-amplitude regions through the external resistance. This accounts for the global electrical feedback (see Section 2) acting equally on all oscillators. Since perfectly homogeneous conditions cannot be achieved in any electrochemical experiments, we consider spatial heterogeneity by assuming that the intrinsic frequencies $\omega(\vec{x})$ vary slightly across the electrode surface. Finally, spatial coupling through ionic and electronic transport must be considered because the silicon electrode is an extended medium; this effect is modeled by a diffusion term. Taken together, we end up with the heterogeneous, globally coupled CGLE:
@@
Here, $\vec{x}=(x_{1},x_{2})$ denotes the spatial coordinates on the rectangular domain $0,L_{1}]\times[0,L_{2}]$. We can represent the complex variable $W$ in polar coordinates with radius $r$ and phase $\phi$:
$W(\vec{x},t)=r(\vec{x},t)\text{e}^{\phi(\vec{x},t)}$ (7)
Next, we discuss the model’s parameters and their relationship to the experiment described above. The shear parameter of the individual oscillators, $C_{2}$, quantifies the amplitude–frequency interaction, whereas the $\omega(\vec{x})$ term specify the natural frequency of each oscillator at position $\vec{x}$. The global coupling parameters are $K$ (diagonal coupling) and $KC_{1}$ (off-diagonal coupling), where $K$ sets the coupling strength and $C_{1}$ modulates the phase-shift introduced by the coupling. Changes in the external resistance modify the potential drop (see equation (1)) and, through this term, affect both the dissipative and reactive components of the effective coupling between oscillators. We therefore assume that the resistance $R$ strongly influences the phase-shift $C_{1}$. Since the microscopic form of the global coupling cannot be determined explicitly, we employ a generic complex coupling term that can account for a broad range of physically plausible global-coupling scenarios $K(1+iC_{1})$. The diffusion strength is given by $\epsilon$. In summary, we can say that the Stuart-Landau oscillator qualitatively represents the local experimental oscillator. A system of Stuart-Landau oscillators interacting globally and through weak diffusion should provide qualitative insight into the emergence of the observed patterns.
Figure 4: Numerically obtained snapshots of the amplitude (first row) and phase (second row) distributions, and frequency distribution (third row) for $K=1.15,C_{2}=2,\epsilon=0.02,\sigma_{0}=0.3$. The $C_{1}$ parameter is adiabatically changed, and the simulation results are plotted for four different values of $C_{1}$. First column $C_{1}=-2.5$, second column $C_{1}=-2.1$, third column $C_{1}=-1.6$, and last column $C_{1}=-1.4$.
We numerically solve Eqs. (5) on a square domain ($L_{1}=L_{2}=10$) with Neumann boundary conditions, discretized on a $100\times 100$ uniform grid. To ensure reproducible frequency realizations, we generate $\omega(\vec{x})$ with Julia’s Mersenne Twister PRNG (seed 2024). The intrinsic frequencies $\omega_{k}$ are drawn from a Gaussian distribution with mean $\mu_{0}$ and standard deviation $\sigma_{0}$; moving to a co-rotating frame with $\mu_{0}$ as its center shifts the distribution to zero. We use adiabatic continuation and integrate system (5) for $T=1000$ time steps. Here, "adiabatic" means that we first run the system for $T$ time steps at $C_{1}=C_{1}^{0}$. We then increase (decrease) $C_{1}$ to $C_{1}^{0}+\Delta C_{1}$ ($C_{1}^{0}-\Delta C_{1}$) and integrate for another T time steps using the final state of the previous run as the initial condition for the next one. The numerical integrations are performed using Julia’s DifferentialEquations.jl ecosystem [Rackauckas.2017]. We use Tsit5 [Tsitouras.2011], a Runge-Kutta-based solver with adaptive timesteps, and save data points at $\Delta t=0.1$. The parameters $K$ and $C_{2}$ are chosen to obtain solutions where phase-locked solutions with fixed amplitude exist [Nakagawa.1993].
Starting from $C_{1}=-1.75$, we vary the coupling parameter in increments of $\Delta C_{1}=\pm 0.05$ until reaching $C_{1}=-1$ and $C_{1}=-3$, respectively. Figure 4 shows the numerical results of the spatial distribution of the amplitude (first row), phase (second row), and frequency (third row) for four different $C_{1}$-values. The amplitude, phase, and frequency are shown in snapshots. All results are recorded after a sufficient evolution time to ensure that transients have decayed.
@@
In Figure 5, we plot the distribution of the amplitudes for the same $C_{1}$-values used in Figure 4. The amplitudes can be extracted from the numerical data using Eq. (7). From the data, it is clear that there are two peaks for all four values of $C_{1}$. Using the Kernel density estimate defined in equation (4), we can locate the mean radius of each mode. The maxima of the distributions, shown as blue dots, are used to quantify the mode states via the ratio $\rho$ defined in Section 2. From the four different histograms, one can clearly observe how oscillators transition from the low-amplitude mode to the high-amplitude mode as $C_{1}$ is increased.
@@
In Figure 6, we plot the fraction of the high-amplitude mode $\rho$ as a function of $C_{1}$ for three different simulation schemes. The cluster ratio $\rho$ has been determined as explained for Figure 3. Since the experimental peaks are broader than the simulated ones, we restrict the counting to oscillators within $\pm 0.15$ of each peak maximum. A unimodal distribution is assigned to $\rho=0$. The blue open circles represent simulation results, where, for $C_{1}=-1.75$, a random initial condition was chosen. For the other values of $C_{1}$, the simulations were continued adiabatically. For the simulations shown in orange symbols, the same initial condition is chosen for each value of $C_{1}$, and the simulation shown in green is identical to orange, but without diffusion. First, consider the adiabatic continuation. Starting at $C_{1}=-3$, the system is in a synchronous solution. At $C_{1}\approx-2.8$, it jumps to a pattern of $\rho\approx 0.06$. In this configuration, many oscillators are in the low-amplitude region, and few are in the high-amplitude region. Increasing $C_{1}$ leads to more oscillators transitioning to the high amplitude region until a critical ratio is reached at $C_{1}\approx-1.1$, at which the system transitions back to a unimodal distribution. For the continuation with a random initial condition, the $C_{1}$-interval where the bimodal pattern appears is narrower. For the no-diffusion series (green), the amplitude distribution is often not clearly bimodal; in those cases, the KDE peak–picking ansatz becomes ill-posed and the inferred $\rho$ fluctuates strongly. Therefore, these apparent “jumps” should be interpreted cautiously, as they largely reflect estimator instability rather than genuine rapid changes of the domain fraction.